AB INITIO CLASSICAL DYNAMICS SIMULATIONS OF CO_{2} LINE-MIXING EFFECTS IN INFRARED BANDS

\underline{JULIEN LAMOUROUX}^{a}, JEAN-MICHEL HARTMANN, HA TRAN, Laboratoire Interuniversitaire des Systèmes Atmosphériques (LISA, CNRS UMR 7583),Université Paris-Est Créteil, Université Paris Diderot, Institut Pierre-Simon Laplace, 94010 Créteil Cedex, France; MAR-CEL SNELS, ISAC-CNR, Via del Fosso del Cavaliere, 100 00133 Rome, Italy; STEFANIA STE-FANI and GIUSEPPE PICCIONI, IAPS-IASF, Via del Fosso del Cavaliere, 100 00133 Rome, Italy.

Ab initio calculations of line-mixing effects in CO_{2} infrared bands are presented and compared with experiments. The predictions were carried using requantized Classical Dynamics Molecular Simulations (rCDMS) based on an approach previously developed and successfully tested for CO_{2} isolated line shapes^{b}. Using classical dynamics equations, the force and torque applied to each molecule by the surrounding molecules (described by an ab initio intermolecular potential) are computed at each time step. This enables, using a requantization procedure, to predict dipole and isotropic polarizability auto-correlation functions whose Fourier-Laplace transforms yield the spectra. The quality of the rCDMS calculations is demonstrated by comparisons with measured spectra in the spectral regions of the 3ν_{3} and 2ν_{1} + 2ν_{2} + ν_{3} Infrared bands^{c}.

aSupport by the French Agence Nationale de la Recherche (ANR) is gratefully acknowledged.

bJ.-M. Hartmann, H. Tran, N. H. Ngo, et al., Phys. Rev. Lett. A \underline{87} (2013), 013403.

cH. Tran, C. Boulet, M. Snels, S. Stefani, J. Quant. Spectrosc. Radiat. Transfer \underline{112} (2011), 925-936.